13. The current spot rate ($/Peso) is .0704. Assume that relative PPP holds, and U.S. inflation is expected to be 3% per year for the next two years while Mexican inflation is expected to be 9% per year over the same period. XYZ Corp of USA has a 20 million peso payable at the end of two years. Calculate the expected amount of dollars needed to make the payment at the end of two years.

14. The current spot rate ($/€) is 1.14. Assume that relative PPP holds between U.S. and Germany, and U.S. inflation is expected to be 3% per year for the next five years while German inflation is expected to be 5% per year over the same period. XYZ Corp of USA has a 10 million Euro receivable at the end of five years. Calculate the expected amount of dollars to be received by XYZ at the end of five years.
Answer #1

13. Exchange rate = Spot Rate*[(1+ r1)/(1+r2)]^n

=  0.0704*[(1+3%)/(1+9%)]^2

= 0.062863

So, 20 mn peso = 20000000*0.062863 = $ 1257257 (Ans)

14. Exchange rate = Spot Rate*[(1+ r1)/(1+r2)]^n

= 1.14*[(1+3%)/(1+5%)]^5

= 1.035487

So, 10 mn Euro = 20000000*1.035487 = $ 10354866 (Ans)